### v3

@@ 7.1.2 Algorithm and Theoretical Bound
Note that with the LP reformulation $V^{\mathrm{Fi}}$ in (43), our Algorithm 1 and Algorithm 2 apply to the episodic setting. The only change we need to make is to replace the matrix $B$ and $C$ with the one described in Section 7.1.1. Denote by $\bm{q}^{1},\dots,\bm{q}^{N}$ the output of Algorithm 2. Then, We define $\bar{\bm{q}}^{N}$ such that $\bar{\bm{q}}^{N}_{\mathcal{I}^{*c}}=0$ and
$\bar{\bm{q}}^{N}_{\mathcal{I}^{*}}=\frac{1}{N}\cdot\sum_{n=1}^{N}\bm{q}^{n}_{\mathcal{I}_{N}}.$
Here, $\mathcal{I}^{*}$ can be replaced by $\mathcal{I}_{N}$, the output of Algorithm 1 with the input sample size $N$. Then, we can construct a policy $\bar{\pi}^{N}$ as follows. For each $h\in[H]$ and $(s,a)\in\mathcal{S}\times\mathcal{A}$, it holds
$P(\bar{\pi}^{N}_{h}(s)=a)=\left\{\begin{aligned} &\frac{\bar{q}^{N}_{h}(s,a)}{\sum_{a^{\prime}\in\mathcal{A}}\bar{q}^{N}_{h}(s,a^{\prime})},&&\text{if\leavevmode\nobreak\ }\sum_{a^{\prime}\in\mathcal{A}}\bar{q}^{N}_{h}(s,a^{\prime})>0\\ &1/|\mathcal{A}|,&&\text{if\leavevmode\nobreak\ }\sum_{a^{\prime}\in\mathcal{A}}\bar{q}^{N}_{h}(s,a^{\prime})=0,\end{aligned}\right.$ (47)
where $P(\bar{\pi}^{N}_{h}(s)=a)$ denotes the probability for the policy $\bar{\pi}$ to take the action $a\in\mathcal{A}$ given the state $s\in\mathcal{S}$ at period $h\in[H]$. Similarly, if $\sum_{a^{\prime}\in\mathcal{A}}\bar{q}^{N}_{h}(s,a^{\prime})=0$, we can take an arbitrary action. Then, we have the following bound regarding the sample complexity of the policy $\bar{\pi}^{N}$, as a corollary of Theorem 6.4.
@@
$O\left(\frac{(|\mathcal{S}|\cdot H+K)^{3}\cdot|\mathcal{S}|\cdot|\mathcal{A}|}{\xi\cdot\sigma^{\prime 3}\cdot\Delta}\cdot\frac{\log^{2}(1/\varepsilon)}{\varepsilon}\right),$
@@
where $\xi=\min_{(s,a)\in\mathcal{I}^{*}}\left\{q^{*}(s,a)\right\}$ and $\bm{q}^{*}$ denotes the optimal soluton to LP (8) corresponding to the optimal basis $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$, we obtain a policy $\bar{\pi}^{N}$ from Algorithm 2 (defined in (47)) such that
$V_{r}(\pi^{*},\mu_{1})-V_{r}(\bar{\pi}^{N},\mu_{1})\leq\varepsilon\text{\leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ }V_{k}(\bar{\pi}^{N},\mu_{1})-\alpha_{k}\leq\varepsilon,\leavevmode\nobreak\ \forall k\in[K].$
7.2 Beyond Generative Model
@@
Note that in the previous sections, we assume the existence of a generative model where we can obtain the sample of reward, cost, and state transition for an arbitrary state and action pair. In this section, we show that our approaches and results can be directly generalized to more involved sampling environments, i.e., the online learning and offline learning environments. We illustrate under the infinite horizon discounted setting described in Section 2. However, our results can also be directly applied to the finite horizon episodic setting described in Section 7.1.
7.2.1 Online Learning
@@
Under the online learning setting, we interact with the true environment to obtain samples. To be specific, suppose that at round $n$ the environment is at state $s_{n}$. Then, in order to obtain the samples, we need to choose an action, denoted by $a_{n}$, and we only obtain the sample of $r(s_{n},a_{n})$, $c_{k}(s_{n},a_{n})$ for each $k\in[K]$, and the state transition $(s_{n},a_{n})\rightarrow s_{n+1}$. This will push the environment to state $s_{n+1}$ and we can only obtain information about state $s_{n+1}$ at the next round. Compared to the generative model described in Section 2 where we can obtain samples for an arbitrary state and action pair, the online learning setting is more restricted in that we can only obtain information about the state we are currently at. The online learning setting is formally described in the following assumption. {assumption} Denote by $\mathcal{M}$ the true environment, with the state space $\mathcal{S}$, action space $\mathcal{A}$, and the transition kernel $P$. Given the current state $s$ and the current action $a$, we obtain a sample of $r(s,a)$, $c_{k}(s,a)$ for each $k\in[K]$, and the state transition $(s,a)\rightarrow s^{\prime}$ following the transition kernel. Then, the state of $\mathcal{M}$ will transit to $s^{\prime}$.
@@
The generalization to the online learning setting is direct. We adopt the idea of off-policy learning, i.e., we use a behavior policy that is different from our policy to generate samples and construct estimates of the underlying CMDP. We then use the obtained samples to form the sample set $\mathcal{F}$ to carry out Algorithm 1 and Algorithm 2. Similar to Li et al. (2021a), we use a stationary policy $\pi_{s}$ as the behavior policy. Given the stationarity of $\pi_{s}$, the sample trajectory can be viewed as a sample path of a time-homogeneous Markov chain over the set of state-action pairs $\mathcal{S}\times\mathcal{A}$. Following Paulin (2015), we make the following uniform ergodicity assumption. {assumption} The Markov chain induced by the stationary behavior policy $\pi_{s}$ is uniformly ergodic. Note that a key property to consider is how often we can obtain a sample of a state-action pair $(s,a)$, which measures the learning efficiency under the stationary behavior policy. Denote by $\mu_{s}$ the stationary distribution over all state-action pairs under the behavioral policy. We then define
@@
Suppose that we obtain $N$ samples under the stationary distribution $\mu_{s}$. Then, following the standard Hoeffding’s inequality, we know that we obtain the number of samples at least $\frac{\mu_{\min}\cdot N}{2}$ for each state-action pair $(s,a)$, with a high probability $1-O(1/N)$. Therefore, we can just pick the number of samples $\frac{\mu_{\min}\cdot N}{2}$ for each state-action pair $(s,a)$ to construct the data set to serve as input to Algorithm 1 and Algorithm 2. Our algorithms will be the same as the ones under the generative model where we query the model for $\frac{\mu_{\min}\cdot N}{2}$ times. We have the following sample complexity bound under the online learning setting as a corollary of Theorem 6.4.
@@
For any $\varepsilon>0$, with a sample complexity bound of
@@
$O\left(\frac{(|\mathcal{S}|+K)^{3}\cdot|\mathcal{S}|\cdot|\mathcal{A}|}{\mu_{\min}\cdot\alpha^{2}\cdot\xi\cdot\sigma(1-\gamma)\cdot\min\{\sigma^{2},(1-\gamma)^{2}\cdot\Delta\}}\cdot\frac{\log^{2}(1/\varepsilon)}{\varepsilon}\right),$
where $\mu_{\min}$ is given in (48), $\xi=\min_{(s,a)\in\mathcal{I}^{*}}\left\{q^{*}(s,a)\right\}$ and $\bm{q}^{*}$ denotes the optimal soluton to LP (8) corresponding to the optimal basis $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$, we obtain a policy $\bar{\pi}^{N}$ from Algorithm 2 such that
$V_{r}(\pi^{*},\mu_{1})-V_{r}(\bar{\pi}^{N},\mu_{1})\leq\varepsilon\text{\leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ }V_{k}(\bar{\pi}^{N},\mu_{1})-\alpha_{k}\leq\varepsilon,\leavevmode\nobreak\ \forall k\in[K].$
7.2.2 Offline Learning
In the offline learning setting, the agent does not interact with the true environment. Instead, the learning task of the agent is conducted using an offline dataset, where it is assumed that the state-action pair follows a reference distribution $\mu_{r}$ (e.g. Rashidinejad et al. (2021), Chen et al. (2022)). Note that in order to implement Algorithm 1 and Algorithm 2, we obtain samples following a uniform distribution $\mu_{u}$ over $\mathcal{S}\times\mathcal{A}$, by querying the generative model. To characterize the distribution shift, we define the following notion of deviation
$D=\max_{(s,a)\in\mathcal{S}\times\mathcal{A}}\left\{\frac{1}{|\mathcal{S}|\cdot|\mathcal{A}|\cdot\mu_{r}(s,a)}\right\}.$ (49)
Then, following the same argument as in the online learning setting, we obtain the following sample complexity bound under the offline learning setting as a corollary of Theorem 6.4.
Corollary 7.4
For any $\varepsilon>0$, with a sample complexity bound of
$O\left(\frac{(|\mathcal{S}|+K)^{3}\cdot D}{\alpha^{2}\cdot\xi\cdot\sigma(1-\gamma)\cdot\min\{\sigma^{2},(1-\gamma)^{2}\cdot\Delta\}}\cdot\frac{\log^{2}(1/\varepsilon)}{\varepsilon}\right),$
@@ 7.2.2 Offline Learning
where $D$ is given in (49), $\xi=\min_{(s,a)\in\mathcal{I}^{*}}\left\{q^{*}(s,a)\right\}$ and $\bm{q}^{*}$ denotes the optimal solution to LP (8) corresponding to the optimal basis $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$, we obtain a policy $\bar{\pi}^{N}$ from Algorithm 2 such that
$V_{r}(\pi^{*},\mu_{1})-V_{r}(\bar{\pi}^{N},\mu_{1})\leq\varepsilon\text{\leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ }V_{k}(\bar{\pi}^{N},\mu_{1})-\alpha_{k}\leq\varepsilon,\leavevmode\nobreak\ \forall k\in[K].$
7.3 Numerical Studies

### v4

@@ 7.1.2 Algorithm and Theoretical Bound
Note that with the LP reformulation $V^{\mathrm{Fi}}$ in (45), our Algorithm 2 and Algorithm 3 apply to the episodic setting. The only change we need to make is to replace the matrix $B$ and $C$ with the one described in Section 7.1.1. Denote by $\bm{q}^{1},\dots,\bm{q}^{N_{2}}$ the output of Algorithm 3. Then, We define $\bar{\bm{q}}^{N_{2}}$ such that $\bar{\bm{q}}^{N_{2}}_{\hat{\mathcal{I}}^{c}}=0$ and
$\bar{\bm{q}}^{N_{2}}_{\hat{\mathcal{I}}}=\frac{1}{N_{2}}\cdot\sum_{n=1}^{N_{2}}\bm{q}^{n}_{\hat{\mathcal{I}}}.$
Here, $\hat{\mathcal{I}}$ refers to the output of Algorithm 2 during the execution of Algorithm 3. Then, we can construct a policy $\bar{\pi}$ as follows. For each $h\in[H]$ and $(s,a)\in\mathcal{S}\times\mathcal{A}$, it holds
[…]
where $P(\bar{\pi}_{h}(s)=a)$ denotes the probability for the policy $\bar{\pi}$ to take the action $a\in\mathcal{A}$ given the state $s\in\mathcal{S}$ at period $h\in[H]$. Similarly, if $\sum_{a^{\prime}\in\mathcal{A}}\bar{q}^{N_{2}}_{h}(s,a^{\prime})=0$, we can take an arbitrary action. Then, we have the following bound regarding the sample complexity of the policy $\bar{\pi}$, as a corollary of Theorem 6.5.
@@
$N=O\left(\frac{H|\mathcal{S}||\mathcal{A}|\cdot\log(1/\varepsilon)}{\min\{\delta_{1}^{2},\delta_{2}^{2},\sigma_{0}^{2}\}}+\frac{m^{7}\cdot\log(1/\xi)}{\xi(\sigma^{*})^{3}(1-\gamma)}\cdot\frac{\log^{2}(1/\varepsilon)}{\varepsilon}\right),$
@@
where $\xi=\min_{h\in[H],(s,a)\in\mathcal{I}^{*}}\left\{q_{h}^{*}(s,a)\right\}$ and $\bm{q}^{*}$ denotes the optimal solution to LP (8) corresponding to the optimal basis $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$, we obtain a policy $\bar{\pi}$ from Algorithm 3 (defined in (49)) such that
$V_{r}(\pi^{*},\mu_{1})-V_{r}(\bar{\pi},\mu_{1})\leq\varepsilon\text{\penalty 10000\ \penalty 10000\ and\penalty 10000\ \penalty 10000\ }V_{k}(\bar{\pi},\mu_{1})-\alpha_{k}\leq\varepsilon,\penalty 10000\ \forall k\in[K].$
7.2 Off-policy Learning Setting.
@@
Note that in the previous sections, we assume the existence of a generative model where we can obtain the sample of reward, cost, and state transition for an arbitrary state and action pair. In this section, we show that our approaches and results can be directly generalized to more involved sampling environments, i.e., the offline learning environments, and we discuss the extension to the online learning setting in the next Section 7.3. We illustrate under the infinite horizon discounted setting described in Section 2. However, our results can also be directly applied to the finite horizon episodic setting described in Section 7.1.
@@
Under the off-policy learning setting, we interact with the true environment to obtain samples by following a behavior policy. Note that the behavior policy is purely for collecting samples and can be different from the target policy. To be specific, suppose that at round $n$ the environment is at state $s_{n}$. Then, in order to obtain the samples, we need to choose an action, denoted by $a_{n}$, and we only obtain the sample of $r(s_{n},a_{n})$, $c_{k}(s_{n},a_{n})$ for each $k\in[K]$, and the state transition $(s_{n},a_{n})\rightarrow s_{n+1}$. This will push the environment to state $s_{n+1}$ and we can only obtain information about state $s_{n+1}$ at the next round. Compared to the generative model described in Section 2 where we can obtain samples for an arbitrary state and action pair, the off-policy learning setting is more restricted in that we can only obtain information about the state we are currently at. The off-policy learning setting is formally described in the following assumption. {assumption} Denote by $\mathcal{M}$ the true environment, with the state space $\mathcal{S}$, action space $\mathcal{A}$, and the transition kernel $P$. Given the current state $s$ and the current action $a$, we obtain a sample of $r(s,a)$, $c_{k}(s,a)$ for each $k\in[K]$, and the state transition $(s,a)\rightarrow s^{\prime}$ following the transition kernel. Then, the state of $\mathcal{M}$ will transit to $s^{\prime}$.
@@
The generalization to the off-policy learning setting is direct as we use a behavior policy that is different from our policy to generate samples and construct estimates of the underlying CMDP. We then use the obtained samples to form the sample set $\mathcal{F}$ to carry out Algorithm 2 and Algorithm 3. Similar to li2021sample, we use a stationary policy $\pi_{s}$ as the behavior policy. Given the stationarity of $\pi_{s}$, the sample trajectory can be viewed as a sample path of a time-homogeneous Markov chain over the set of state-action pairs $\mathcal{S}\times\mathcal{A}$. Following paulin2015concentration, we make the following uniform ergodicity assumption. {assumption} The Markov chain induced by the stationary behavior policy $\pi_{s}$ is uniformly ergodic. Note that a key property to consider is how often we can obtain a sample of a state-action pair $(s,a)$, which measures the learning efficiency under the stationary behavior policy. Denote by $\mu_{s}$ the stationary distribution over all state-action pairs under the behavioral policy. We then define
@@
Suppose that we obtain $N$ samples under the stationary distribution $\mu_{s}$. Then, following the standard Hoeffding’s inequality, we know that we obtain the number of samples at least $\frac{\mu_{\min}\cdot N_{1}}{2}$ for each state-action pair $(s,a)$, with a high probability $1-O(1/N_{1})$. Therefore, we can just pick the number of samples $\frac{\mu_{\min}\cdot N_{1}}{2}$ for each state-action pair $(s,a)$ to construct the data set to serve as input to Algorithm 2 in the execution of Algorithm 3. For the resolving procedure in Algorithm 3 and for each round of iteration, we collect samples using the behavior policy until we collect one sample for each $(s,a)$. We then have the following sample complexity bound under the off-policy learning setting, which generalizes the analysis in Theorem 6.5.
@@
For any $\epsilon>0$, with a sample complexity bound of
@@
$N=O\left(\frac{\log(1/\varepsilon)}{\mu_{\min}\cdot\min\{\delta_{1}^{2},\delta_{2}^{2},\sigma_{0}^{2}\}}+\frac{m^{5}\cdot\log(1/\xi)}{\mu_{\min}\cdot\xi(\sigma^{*})^{3}(1-\gamma)}\cdot\frac{\log^{2}(1/\varepsilon)}{\varepsilon}\right),$
@@
where $\mu_{\min}$ is given in (50) and $\xi=\min_{(s,a)\in\mathcal{I}^{*}}\left\{q^{*}(s,a)\right\}$ and $\bm{q}^{*}$ denotes the optimal solution to LP (8) corresponding to the optimal basis $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$, we obtain a policy $\bar{\pi}$ from Algorithm 3 such that
$V_{r}(\pi^{*},\mu_{1})-V_{r}(\bar{\pi},\mu_{1})\leq\varepsilon\text{\penalty 10000\ \penalty 10000\ and\penalty 10000\ \penalty 10000\ }V_{k}(\bar{\pi},\mu_{1})-\alpha_{k}\leq\varepsilon,\penalty 10000\ \forall k\in[K].$
7.3 On-policy Learning Setting.
Different from the off-policy learning setting, under the on-policy learning setting, we interact with the true environment to obtain samples by following the target policy. To be specific, suppose that at round $n$ the environment is at state $s_{n}$. Then, denoting by $\bm{q}_{t}$ the occupancy measure computed by our algorithm, in order to obtain the samples, we need to choose an action following the occupancy measure $\bm{q}_{t}$, denoted by $a_{n}$, and we only obtain the sample of $r(s_{n},a_{n})$, $c_{k}(s_{n},a_{n})$ for each $k\in[K]$, and the state transition $(s_{n},a_{n})\rightarrow s_{n+1}$. This will push the environment to state $s_{n+1}$ and we can only obtain information about state $s_{n+1}$ at the next round. Compared to the generative model described in Section 2 where we can obtain samples for an arbitrary state and action pair, the on-policy learning setting is more restricted in that we can only obtain information about the state we are currently at. The on-policy learning setting is even more restrictive than the off-policy learning setting described in Section 7.2 in that we cannot adopt a behavior policy to collect samples, instead, our algorithm needs to satisfy certain properties such that on one hand, we can approximate the optimal policy with instance-dependent guarantees, and on the other hand, we can collect enough samples over the state-action pairs we are concerned about.
Our algorithms can be described as follows. There are two phases of our algorithms. In the first phase, we still use Algorithm 2 to identify the optimal basis $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$. Note that the policy will be a stationary policy, and following the same procedure of Theorem 6.1, we can show that Algorithm 2 enables us to identify one optimal basis with a high probability within a finite time steps scaling as $O(\log(1/\epsilon)/\Delta)$. Then, in the second phase, we apply the resolving procedure in Algorithm 3 during the remaining time periods. Denoting by $\bm{q}^{*}$ the optimal occupancy measure obtained from solving the LP (8) corresponding to the optimal basis $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$, we can show that during the resolving procedure after a warm start in the first phase, the $\bm{q}^{n}$ we obtain is always positively supported on $\mathcal{I}^{*}$. Therefore, if we follow the policy induced by $\bm{q}^{n}$ at each iteration $n$ for a constant number of times, we can actually show that we are able to collect at least one sample over the costs and the state transition associated with each $(s,a)\in\mathcal{I}^{*}$, with a high probability, which is already enough to estimate the matrix $A^{*}=\begin{bmatrix}C(\mathcal{J}_{1}^{*},\mathcal{I}^{*})\\ B(\mathcal{J}_{2}^{*},\mathcal{I}^{*})\end{bmatrix}$ more and more accurately during the resolving procedure. This result shows that our resolving procedure would automatically guarantee exploitation and exploration at the same time, and there is no need to do additional exploration in our algorithmic design. The formal algorithm for the online setting is presented in Algorithm 4.
Algorithm 4 The Adaptive-resolving Algorithm for On-policy Learning
1:Input: the number of rounds $N_{1}$ for carrying out Algorithm 2, the number of rounds $N_{2}$ for resolving, the limit of the number of steps $N_{3}$ for sampling, an estimate $\xi^{\prime}$ such that $\frac{\xi}{4}\leq\xi^{\prime}\leq\xi/2$.
2:Follow a stationary policy by $N_{1}$ periods and obtain the dataset $\mathcal{F}_{N_{1}}$.
3:Obtain the output $\hat{\mathcal{I}}$ and $\hat{\mathcal{J}}$ from Algorithm 2 with input $\mathcal{F}_{N_{1}}$ and $\epsilon=O(\varepsilon^{2}/(K\cdot|\mathcal{S}||\mathcal{A}|))$.
4:Let $\hat{\mathcal{J}}_{1}=\hat{\mathcal{J}}\cap[K]$ and $\hat{\mathcal{J}}_{2}=\hat{\mathcal{J}}\cap\mathcal{S}$.
5:Initialize $\mathcal{H}_{1}=\emptyset$, $\bm{\alpha}^{1}=N_{2}\cdot\bm{\alpha}$, $\bm{\mu}^{1}=N_{2}\cdot\bm{\mu}$, and a state $s$ such that $(s,a)\in\hat{\mathcal{I}}$ for an $a$.
6:for $n=1,\dots,N_{2}$ do
[…]
8:  Construct a solution $\tilde{\bm{q}}^{n}$ such that $\tilde{\bm{q}}^{n}_{\hat{\mathcal{I}}^{c}}=0$ and $\tilde{\bm{q}}^{*}_{\hat{\mathcal{I}}}$ is the solution to
$\begin{bmatrix}\bar{C}^{n}(\hat{\mathcal{J}}_{1},\hat{\mathcal{I}})\\ \bar{B}^{n}(\hat{\mathcal{J}}_{2},\hat{\mathcal{I}})\end{bmatrix}\cdot\tilde{\bm{q}}^{n}_{\hat{\mathcal{I}}}=\begin{bmatrix}\frac{\bm{\alpha}^{n}_{\hat{\mathcal{J}}_{1}}}{N_{2}-n+1}\\ \frac{\bm{\mu}^{n}_{\hat{\mathcal{J}}_{2}}}{N_{2}-n+1}\end{bmatrix}$. (51)
9:  Project $\tilde{\bm{q}}^{n}$ to the set $\{\bm{q}:\|\bm{q}\|_{1}\leq 2,\bm{q}_{\hat{\mathcal{I}}^{c}}=0,\bm{q}_{\hat{\mathcal{I}}}\geq\xi^{\prime}\}$ to obtain $\bm{q}^{n}$.
10:  Construct a policy $\bar{\pi}^{n}$ such that
[…]
11:  Follow the policy $\bar{\pi}^{n}$ until we visit each $(s,a)\in\hat{\mathcal{I}}$ or number of steps reaches the limit $N_{3}$.
12:  Collect sample of $r^{n}(s,a)$, $c^{n}_{k}(s,a)$ for each $k\in\hat{\mathcal{J}}_{1}\subset[K]$, and $s^{n}(s,a)$, for $(s,a)\in\hat{\mathcal{I}}$.
13:  Update $\mathcal{H}_{n+1}=\mathcal{H}_{n}\cup\left\{r^{n}(s,a),c^{n}_{k}(s,a),s^{n}(s,a),\forall(s,a)\in\hat{\mathcal{I}},\forall k\in\hat{\mathcal{J}}_{1}\right\}$.
14:  Denote by $\bm{c}^{n}(s,a)=(c^{n}_{k}(s,a))_{\forall k\in\hat{\mathcal{J}}_{1}}$ and do the update:
$\bm{\alpha}^{n+1}_{\hat{\mathcal{J}}_{1}}=\bm{\alpha}^{n}_{\hat{\mathcal{J}}_{1}}-\sum_{(s,a)\in\hat{\mathcal{I}}}\bm{c}^{n}(s,a)\cdot q^{n}(s,a).$ (53)
15:  Do the update:
$\bm{\mu}^{n+1}_{s^{\prime}}=\bm{\mu}^{n}_{s^{\prime}}-\sum_{(s,a)\in\hat{\mathcal{I}}}q^{n}(s,a)\cdot(\delta_{s^{\prime},s}-\gamma\mathbbm{1}_{\{s^{\prime}=s^{n}(s,a)\}}),\penalty 10000\ \penalty 10000\ \penalty 10000\ \forall s^{\prime}\in\hat{\mathcal{J}}_{2}.$ (54)
16:end for
17:Output: $\bm{q}^{1},\dots,\bm{q}^{N_{2}}$.
We now conduct the theoretical analysis of our algorithm. The analysis of the Algorithm 2 in Algorithm 4 is the same as Theorem 6.1, and we can show that it holds $\hat{\mathcal{I}}$$\mathcal{I}^{*}$ and $\hat{\mathcal{J}}=\mathcal{J}^{*}$ with a high probability. In order to analyze the resolving procedure, a key point is to characterize the distance between $\bm{q}^{n}$ and $\bm{q}^{*}$, and show that following the policy $\bar{\pi}^{n}$ induced by $\bm{q}^{n}$ would guarantee automatic exploration over the entire set $\mathcal{I}^{*}$ for collecting samples over costs and state transition. In fact, denoting by $\xi=\min_{(s,a)\in\mathcal{I}^{*}}\left\{q^{*}(s,a)\right\}$, we can show that $\|\bm{q}^{n}-\bm{q}^{*}\|_{\infty}\leq\frac{\xi}{2}$. This would further imply that the induced policy $\pi^{n}$ guarantees that $P(\pi^{n}(s)=a)\geq\frac{\xi}{4}$ for every $(s,a)\in\mathcal{I}^{*}$ and hence each $(s,a)\in\mathcal{I}^{*}$ can be fully explored. We summarize the above arguments in the following lemma.
Lemma 7.4
Suppose that the sets $\hat{\mathcal{I}}$ and $\hat{\mathcal{J}}$ obtained in step 3 of Algorithm 4 satisfy that $\hat{\mathcal{I}}$$\mathcal{I}^{*}$ and $\hat{\mathcal{J}}=\mathcal{J}^{*}$. Then, when $N_{4}\leq n\leq\tau_{0}$, for a constant $N_{4}$ and a stopping time $\tau_{0}$ defined as
$\tau_{0}=\min_{n\in[N_{2}]}\{\left\|\tilde{\bm{\alpha}}_{\mathcal{J}^{*}_{1}}(n)-\bm{\alpha}_{\mathcal{J}^{*}_{1}}\right\|_{\infty}>\nu\text{\penalty 10000\ or\penalty 10000\ }\left\|\tilde{\bm{\mu}}_{\mathcal{J}^{*}_{2}}(n)-\bm{\mu}_{\mathcal{J}^{*}_{2}}\right\|_{\infty}>\nu\}$ (55)
for a constant $\nu$, with a probability at least $1-\epsilon$, it holds that
$\|\bm{q}^{n}-\bm{q}^{*}\|_{\infty}\leq\frac{\xi}{2}$
where $\bm{q}^{n}$ denotes the solution obtained from the resolving procedure of Algorithm 4 and $\bm{q}^{*}$ denotes the optimal solution to LP (8) corresponding to the basis $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$.
In order to further bound the sample complexity, we denote by $\Pi$ the set of policies $\pi$ with the corresponding occupancy measure supported on $\mathcal{I}^{*}$ and it holds $P(\pi(s)=a)\geq\xi^{\prime}$ for every $(s,a)\in\mathcal{I}^{*}$. We also denote by $t_{\mathrm{mix}}(\pi)$ the mixing time of the Markov chain corresponding to the policy $\pi$. We then define
$t_{\mathrm{mix}}=\max_{\pi\in\Pi}\{t_{\mathrm{mix}}(\pi)\}.$ (56)
Denoting by $\bm{q}^{\pi}$ the occupancy measure corresponding to a policy $\pi\in\Pi$, we define
$\xi_{\mathrm{small}}=\min_{\pi\in\Pi}\left\{\min_{(s,a)\in\mathcal{I}^{*}}\{q^{\pi}(s,a)\}\right\}$ (57)
The largest mixing time $t_{\mathrm{mix}}$ and the smallest occupancy measure $\xi_{\mathrm{small}}$ will be used to bound by how many samples we are guaranteed to obtain at least one sample for each $(s,a)\in\mathcal{I}^{*}$ if we are following a policy $\bar{\pi}^{n}$ induced by $\bm{q}^{n}$ as defined in (52). We assume that every Markov chain induced by any $\pi\in\Pi$ is ergodic. We finally use the output $\bm{q}^{1},\dots,\bm{q}^{N_{2}}$ to construct the policy $\bar{\pi}$ by
[…]
where $\bar{\bm{q}}^{N_{2}}=\frac{1}{N_{2}}\cdot\sum_{n=1}^{N_{2}}\bm{q}^{n}$. We have the following sample complexity bound over the on-policy learning setting of our Algorithm 4.
Theorem 7.5
With a sample complexity bound of
$N=O\left(\frac{\log(1/\varepsilon)}{\mu_{\min}\cdot\min\{\delta_{1}^{2},\delta_{2}^{2},\sigma_{0}^{2}\}}+\frac{t_{\mathrm{mix}}\cdot m^{5}\cdot\log(1/\xi)}{\xi_{\mathrm{small}}\cdot(\xi\sigma^{*})^{3}(1-\gamma)}\cdot\frac{\log^{2}(1/\varepsilon)}{\varepsilon}\right),$
where $\mu_{\min}$ is defined in (50), $t_{\mathrm{mix}}$ is defined in (56), and $\xi_{\mathrm{small}}$ is defined in (57), and we obtain a policy $\bar{\pi}$ from Algorithm 3 (defined in (58)) such that
$V_{r}(\pi^{*},\mu_{1})-V_{r}(\bar{\pi},\mu_{1})\leq\varepsilon\text{\penalty 10000\ \penalty 10000\ and\penalty 10000\ \penalty 10000\ }V_{k}(\bar{\pi},\mu_{1})-\alpha_{k}\leq\varepsilon,\penalty 10000\ \forall k\in[K].$
Remark 7.6
Note that in our Algorithm 4, we need to specify the number of rounds $N_{1}$ for carrying out Algorithm 2, the number of rounds $N_{2}$ for resolving, and the limit of the number of periods $N_{3}$ for sampling in each iteration of the resolving procedure. The value of $N_{1}$, $N_{2}$, and $N_{3}$ depends on some problem parameters which could be unknown to the decision maker. However, on one hand, we can estimate these problem parameters up to a constant level of accuracy with a high probability, which would not change the order of our sample complexity bound. On the other hand, we can adopt a standard doubling trick to avoid setting these values explicitly. For example, in order to avoid setting the value of $N_{3}$ explicitly, in each iteration of the resolving procedure, we can initialize with some $N_{3}$ and exponentially increase the value of $N_{3}$ until we have at least one sample of $(s,a)\in\hat{\mathcal{I}}$. The final value of $N_{3}$ will be bounded by a constant with a high probability.
8 Concluding Remarks
In this work, we present a new algorithm that achieves logarithmic regrets or $\tilde{O}(1/\epsilon)$ sample complexity bound for the CMDP problems, for the first time in the literature. To achieve these improvements, we develop a new algorithmic framework to analyze the CMDP problems, motivated by the online LP literature. We consider the LP reformulation of the CMDP problems and we solve the LP in an online manner. Besides introducing the online LP framework, our method is new even compared to the online LP literature, as we do not require the non-degeneracy assumption to derive the logarithmic regret. In our research, we adopt the basic tabular setting and the infinite-horizon discounted setting to derive our results. We further extend to the finite-horizon episodic setting, off-policy learning setting, and the on-policy learning setting. There are also multiple ways to further extend our approaches. For example, one may incorporate the function approximation to deal with the continuous state-action space setting, or consider the setting where the state and the reward are not directly observed. There could also be other ways to characterize the instance hardness other than the LP-based corner point approach. We leave deriving instance-dependent guarantees for more involved settings in CMDP problems for future research.
[…]
{APPENDICES}
9 Numerical Studies